OpenStudy (anonymous):

can anyone give me some examples of problems like " whats is the solution set for (x=1)^2=49?

OpenStudy (radar):

Have you stated the example correctly?

OpenStudy (anonymous):

dang it. no...it was (x+1)^2=49

OpenStudy (anonymous):

(x+1)^2=7^2 so (x+1)=plus minus (7) x=plus minus 6

OpenStudy (radar):

O.K look like 6 would do it so would -8. Lets work it out. \[(x+1)^{2}= x ^{2}+2x+1=49\] \[x ^{2}+2x-48=0\] \[(x+8)(x-6)=0\]

OpenStudy (anonymous):

umm..the answer choices are in format of ( -48,48) and that kind of thing

OpenStudy (radar):

(x-6)=0, x=6 (x+8)=0, x=-8

OpenStudy (radar):

The solution set is x= 6, -8

OpenStudy (radar):

What exactly is asked for the example?

OpenStudy (radar):

Are you working on the "completing the square" method, or factoring or quadratic? for solutions??

OpenStudy (anonymous):

just solution sets. its part of the algebra 2 EOC, so i want to learn how to work them out

OpenStudy (radar):

method I used is called the "factoring method" The quadratic equation was manipulated to equal 0 and then factored. (x+8)(x-6) =0 The reason this works, one or both of the factors has to equal 0 for the results to equal 0. Like, ab=0, then a or b or both equal 0.

OpenStudy (anonymous):

um...im actually not that good at math, just found this site..wish i had found it sooner though. i could have asked for help on my last project